### v1

@@ 4 Resource Scheduling in Venn
In this section, we first introduce the problem statement (§4.1). We then develop our scheduling algorithm in two steps: establishing the job scheduling order to minimize the scheduling delay (§4.2) and determining the device-to-job matching to co-optimize the response collection time (§4.3). Then, we discuss additional enhancements for real-world deployments (§4.4).
@@ 4.1 Problem Statement
Given a collection of FL jobs—along with their device requirements and resource demands—and a set of heterogeneous devices that arrive and depart over time, Venn should efficiently assign devices to FL jobs in order to reduce the average job completion time (JCT). The scheduling problem can be mathematically modeled as a multi-commodity flow (MCF) problem with integer constraints, where each FL job is modeled as a distinct commodity and each device serves as an intermediate vertex between the source and sink of its corresponding eligible FL job. Then the goal of this integer MCF problem is to minimize the average JCT of jobs, which is known to be NP-hard (Even et al., 1975). Even for its linear approximation, the time complexity is exacerbated by the planetary scale of devices involved and diverse device requirements from jobs, making existing solutions computationally infeasible. We now formally define such resource scheduling problem.
@@
Assume we have $m$ jobs $\mathbb{J}=\{J_{1},J_{2},\allowbreak...,J_{m}\}$ with their resource demands $\mathbb{D}=\{D_{1},D_{2},\allowbreak...,D_{m}\}$. Let $\mathbb{S}=S_{1}\cup S_{2}\cup...\cup S_{n}$ be the available device set, where $S_{k}$ is the eligible device subset in $\mathbb{S}$ that satisfies the device specification of job $J_{i}$, i.e., $f(J_{i})=S_{k}$. The goal is to match each checked-in device $s\in S_{k}$ with one job $J_{i}$, where $f(J_{i})=S_{k}$, $\forall s\in\mathbb{S}$, in order to minimize average JCT, which consists of scheduling delay and response collection time.
@@
Jointly optimizing scheduling delay and response collection time is non-trivial. Intuitively, to reduce the average scheduling delay, it is desirable to promptly assign each device upon check-in to one eligible job. On the other hand, since the response collection time is usually determined by the final responding device among the target number of participants, it can be minimized by mitigating stragglers and cherry picking ideal devices to jobs. However, such an approach takes a longer time to acquire a certain number of qualified devices, leading to longer scheduling delay.
At its core, Venn aims to find a sweet spot in the trade-off in order to optimize average JCT. Specifically, we decouple the FL resource allocation problem as following two questions:
@@
    How to decide the FL job scheduling order in order to minimize the average scheduling delay? (§4.2)
@@
    How to match devices with FL jobs in order to minimize average response collection time while also reducing average JCT? (§4.3)
@@
1:function Venn-Sched( Job Groups $\mathbb{G}$, Devices $\mathbb{S}$ )
$\triangleright$    Sort within job group (§4.2.1).
@@
$\triangleright$    Generate initial allocation $S_{j}^{\prime}$ for each group $G_{j}$.
4:     $S=\cup_{j=1}^{n}S_{j}$
5:     Sort $G_{j}$ by $|S_{j}|$ in ascending order, $\forall G_{j}\in\mathbb{G}$
6:     for $G_{j}$ in $G$ do
7:         $S_{j}^{\prime}=S\cap S_{j},S=S\setminus S_{j}^{\prime}$
$\triangleright$    Allocate resource $S_{j}^{\prime}$ for each group $G_{j}$.
8:     Sort $G_{j}$ by $|S_{j}|$ in descending order, $\forall G_{j}\in\mathbb{G}$
9:     for $G_{j}$ in $G$ do
10:         if  $|S_{j}^{\prime}|>0$
11:              for $G_{k}\in\mathbb{G}:|S_{k}|<|S_{j}|,S_{k}\cap S_{j}\neq\emptyset$ do
12:                  $m_{j}^{\prime},m_{k}^{\prime}$= get-queue-len()
13:                  if  $\frac{m_{j}^{\prime}}{|S_{j}^{\prime}|}>\frac{m_{k}^{\prime}}{|S_{k}|}$
14:                       $S_{j}^{\prime}=S_{j}^{\prime}\cup(S_{j}\cap S_{k})$
15:                       $S_{k}^{\prime}=S_{k}^{\prime}-S_{j}^{\prime}$
16:                  else
17:                       break
18:     return $\{G_{j}[0],S_{j}^{\prime}\},\forall j\in[1,n]$
@@ 4.2 Intersection Resource Scheduling (IRS)
We now tackle the first question, minimizing the average scheduling delay. Directly matching jobs to devices can be computationally expensive, especially when dealing with the immense scale of devices and jobs. The challenges posed by this problem are not solely due to its scale either; they are also compounded by the diverse resource requirements of FL jobs. These varying requirements introduce intricate resource contention patterns, further complicating the scheduling process.
To this end, Venn introduces the Intersection Resource Scheduling (IRS) problem to account for this resource contention. Basically, each FL job $J_{i}\in\mathbb{J}$ may compete for a subset of devices $S_{k}\in\mathbb{S}$, denoted as $f(J_{i})=S_{k}$, where these resource subsets can exhibit relationships that are inclusive, overlapping, or nested. We created an integer linear programming (ILP) formulation to optimally allocate resources to minimize scheduling delay (Appendix A) and propose a heuristic to tackle the problem.
@@
With the objective of determining a job scheduling order, Venn first groups jobs $\mathbb{J}$ into Resource-Homogeneous Job Groups $\mathbb{G}=\{G_{1},G_{2},...,G_{n}\}$ by their resource requirements, where each job group $G_{j}=\{J_{i}|f(J_{i})=S_{j},\forall J_{i}\in\mathbb{J}\}_{i=1}^{m_{j}}$ contains all jobs with the same resource requirement. Venn addresses the problem using a two-step approach, with each step occurring at a different level of scheduling granularity, as outlined in Algorithm 1: (i) Determining the job order within a job group to optimize local resource scheduling (§4.2.1). (ii) Then deciding how to merge the job order across job groups to ensure global scheduling efficiency (§4.2.2). Venn invokes Algorithm 1 on job’s request arrival and completion. By breaking down the overall problem into two steps, we further reduce the problem’s complexity without affecting the scheduling efficiency. We provide theoretical insights to show the effectiveness of this problem decomposition in Appendix B.
4.2.1 Intra Job Group Scheduling
@@
4.2.2 Inter Job Group Scheduling
@@
An effective scheduler should recognize and adapt to the resource contention patterns among jobs. At a high-level, it should prioritize jobs with scarce resources, i.e., those with stringent device requirements and limited eligible resource options, to avoid being blocked by jobs with more abundant resources. Additionally, when a particular resource type is in high demand, the scheduler should judiciously allocate intersected resources to the job group with a longer queue, as this group contributes more significantly to the average scheduling delay.
@@
    Amount of eligible resources allocated: the job group with smaller amount of eligible resources may have longer scheduling delay under the same condition.
@@
Algorithm 1 outlines the steps Venn takes to allocate the current resource across job groups. First, Venn initializes resource allocation among job groups by starting to allocate resources to job group with most scarce resources (line 5). This results in an initial allocation plan with no resource sharing across job groups (lines 47), setting the stage for subsequent cross-group allocations.
To determine how to allocate intersected resource across job groups, Venn greedily evaluates whether a job group with more abundant resources should acquire intersected resources from groups with scarcer resources with the objective of minimizing the average scheduling delay. This evaluation starts with the job group possessing the most abundant resources (line 8). If the resources allocated to this group remain unclaimed by other groups (line 10), Venn will decide how much resources from subsequent job groups should be allocated to it. Specifically, Venn prioritizes job groups with longer queue lengths and fewer allocated resources, guided by a ratio that balances the number of affected jobs against the amount of allocated resources (line 13). If this ratio $\frac{m_{j}^{\prime}}{|S_{j}^{\prime}|}$ is larger than the one for the target resource group $\frac{m_{k}^{\prime}}{|S_{k}^{\prime}|}$, Venn reallocates resources accordingly (lines 1415). Otherwise, the algorithm ceases to allocate additional resources to the current job group from remaining groups (line 17). The reason is that if this job group needs more resources, it should first take the resources from job groups with relatively abundant resources.
Function get-queue-len() (line 12) would return the number of jobs $m^{\prime}$ whose JCT would be delayed by potential group prioritization. For example, the affected queue length $m^{\prime}$ may contain jobs from other job groups that have been deprioritized previously. An easier way to approximate $m^{\prime}$ is to use the group queue length. If intersected resources have been decided to be allocated to job group $G_{j}$ over $G_{k}$, Venn accumulates and updates their current resource allocation $S_{j}^{\prime},S_{k}^{\prime}$ and queuing length $m_{j}^{\prime}$ (line 1415).
The time complexity of Algorithm 1 is $\max(O(m\log m),O(n^{2}))$, where $m$ is the number of ongoing jobs and $n$ is the number of job groups. To formalize the effectiveness of Venn’s approach, we illustrate the theoretical insight behind the scheduling algorithm in Appendix C.
@@ 4.3 Device Matching
Response collection time is usually determined by the last successfully responding devices. Hence, it can be reduced by allocating devices with similar higher capacity to the FL job. Meanwhile, these high-end devices have lower probability to fail due to their quick task execution.
Figure 7: Visualize tier-based device-to-job matching condition.
However, as mentioned in Section 4.1, there is trade-off between scheduling delay and response collection time. Intuitively, with limited device influx, priority should be given to minimizing the scheduling delay, which dominates the average JCT. On the other hand, with sufficient device influx to fulfill a job request within a short period, we should consider minimizing the response collection time while obeying the scheduling order given by Section 4.2.
@@
1:$S=\{S^{1},S^{2},...,S^{V}\}$triangleright Evenly partition devices
2:$g_{v}=\frac{t^{v}}{t^{0}},\forall v\in[1,V]$triangleright Response time speed up
3:Input: $J_{i}$, $S_{j}^{\prime}\in{Venn}-SCHED(\mathbb{G},\mathbb{S})$
4:function Venn-Match( Job $J_{i}$, Resource $S_{j}^{\prime}$ )
5:     $c_{i}=\frac{t_{response}}{t_{schedule}}$
$\triangleright$     Assign tiers in a rotating manner.
6:     u = randint(0, V)
$\triangleright$     Decide whether to trigger tier-based matching
7:     if $V+g_{u}*c_{i}<c_{i}+1$
$\triangleright$         Assign tier $u$ devices to $J_{i}$
8:         Update $S_{j}^{\prime}=S_{j}^{\prime}\cap S^{u}$
9:     return $\{J_{i},S_{j}^{\prime}\}$
@@
For each served job request, Venn randomly selects a device tier, denoted as $S^{u}$, to the job (Algorithm 2 line 6). This randomized tier assignment aims to expose each FL job to a diverse set of devices, rather than confining them to high-end devices. Given that the response collection time is determined by the slowest responding participant, tier-based assignment does not adversely affect this metric.
@@
To determine the response time speed-up factor $g$ for tier-based matching, we note that the device response time distribution adheres to a log-normal distribution (Wang et al., 2023). We use the 95th percentile as the statistical tail latency to account for the overall response collection time, thereby excluding failures and stragglers. Venn profiles and estimates the response collection time for each device tier $v\in[1,..,V]$ and subsequently computes the speed-up factor $g_{v}=\frac{t^{v}}{t^{0}}$ relative to a non-tiered scenario (line 2).
@@ 4.4 Enhancements
Dynamic Resource Supply
As shown in Figure 1(a), the total available FL resources change significantly over time. To address this, Venn initially records each device eligibility through a time-series database. This database is then queried for resource eligibility distribution from a past window, which is sourced from public device check-in information. However, relying solely on momentary eligible resource rates for input into the scheduling algorithm is inaccurate. This is primarily due to the varying resource arrival patterns, as demonstrated in Figure 1(a). Given that FL jobs often span multiple days and resource availability typically follows a diurnal pattern, a more accurate approach is to use the average eligible resource rate over a 24-hour period as a representative metric for each job’s eligible resources. As a result, the scheduler can become both farsighted and robust, effectively accommodating the dynamic nature of resource availability.
Starvation Prevention
Our heuristic can lead to larger FL jobs being starved due to the preference given to smaller jobs. This is not acceptable especially when the jobs are initiated by different FL developers who require performance guarantees. Venn grants fairness to jobs to avoid such starvation. Specifically, our goal is to guarantee that the scheduling latency of a job $J_{i}$ is no worse than fair sharing, which is defined as $T_{i}=M*sd_{i}$, where $M$ is the number of simultaneous FL jobs and $sd_{i}$ represents the JCT without contention. Then, we adjust each job demand to be $d_{i}^{\prime}=d_{i}\times(\frac{t_{i}}{T_{i}})^{\epsilon}$ to ensure fairness within a job group, and adjust each group queue length $q_{j}^{\prime}=q_{j}\times(\frac{\sum_{J_{i}\in G_{j}}T_{i}}{\sum_{J_{i}\in G_{j}}t_{i}})^{\epsilon}$ to ensure fairness across job groups. $t_{i}$ is the time usage of job $J_{i}$ at the moment and $\epsilon\in[0,\infty)$ is a fairness control knob. When $\epsilon=0$, the algorithm is identical to the one in Section 4.2. As $\epsilon\rightarrow\infty$, the fairness multiplier dominates the scheduling, resulting in maximum fairness. We show that Venn improves JCT over its counterparts with our starvation design (§5.5).

### v2

@@ 4 Resource Scheduling in Venn
In this section, we outline the resource scheduling algorithm in Venn, starting with defining the scheduling problem in CL (§4.1). Next, we present our scheduling algorithm in two parts: determining the job scheduling order to minimize scheduling delay (§4.2) and matching devices to jobs to optimize response collection time (§4.3). Finally, we describe enhancements for real-world deployment (§4.4).
@@ 4.1 Problem Statement
Given a collection of CL jobs—along with their device requirements and resource demands—and a set of heterogeneous devices that arrive and depart over time, Venn should efficiently assign devices to CL jobs in order to reduce the average job completion time (JCT). The scheduling problem can be mathematically modeled as a multi-commodity flow (MCF) problem with integer constraints, where each CL job is modeled as a distinct commodity and each device serves as an intermediate vertex between the source and sink of its corresponding eligible CL job. Then the goal of this integer MCF problem is to minimize the average JCT of jobs, which is known to be NP-hard (Even et al., 1975). Even for its linear approximation, the time complexity is exacerbated by the planetary scale of devices involved and diverse device requirements from jobs, making existing solutions computationally infeasible.
@@
Let $\mathbb{J}={J_{1},J_{2},\ldots,J_{m}}$ be the set of jobs, with resource demands $\mathbb{D}={D_{1},D_{2},\ldots,D_{m}}$, where $D_{i}$ is the number of devices required by job $J_{i}$. Let $\mathbb{S}=S_{1}\cup S_{2}\cup\cdots\cup S_{n}$ be the set of available devices, where $S_{k}$ is the subset of devices eligible for jobs with specific requirements (e.g., hardware capacity), and $f(J_{i})=S_{k}$ maps job $J_{i}$ to its eligible device subset. The goal is to match each checked-in device $s\in S_{k}$ with one job $J_{i}$, where $f(J_{i})=S_{k}$, $\forall s\in\mathbb{S}$, in order to minimize average JCT, which consists of scheduling delay and response collection time.
@@
Minimizing scheduling delay often conflicts with minimizing response collection time. Assigning devices quickly reduces delay but may involve slower devices, increasing response time, since the response collection time is usually determined by the final responding device among the target number of participants. Waiting for faster devices, however, reduces response time at the cost of higher delay.
At its core, Venn aims to find a sweet spot in the trade-off in order to optimize average JCT. Specifically, we decouple the CL resource allocation problem as following two questions:
@@
    How should jobs be ordered to minimize scheduling delay? (§4.2)
@@
    How should devices be matched to jobs to minimize response collection time while optimizing JCT? (§4.3)
@@
1:  Input: Job Groups $\mathbb{G}$, Devices $\mathbb{S}$ FunctionVenn-SchedJob Groups $\mathbb{G}$, Devices $\mathbb{S}$triangleright Sort within job group (§4.2.1)
@@
[…]
5:  $S=\cup_{j=1}^{n}S_{j}$
6:  Sort $G_{j}$ by $|S_{j}|$ in ascending order, $\forall G_{j}\in\mathbb{G}$
7:  for $G_{j}$ in $G$ do
8:     $S_{j}^{\prime}=S\cap S_{j},S=S\setminus S_{j}^{\prime}$
[…]
10:  Sort $G_{j}$ by $|S_{j}|$ in descending order, $\forall G_{j}\in\mathbb{G}$
11:  for $G_{j}$ in $G$ do
12:     if $|S_{j}^{\prime}|>0$ then
13:        for $G_{k}\in\mathbb{G}:|S_{k}|<|S_{j}|,S_{k}\cap S_{j}\neq\emptyset$ do
14:           $m_{j}^{\prime},m_{k}^{\prime}$ = get-queue-len()
15:           if $\frac{m_{j}^{\prime}}{|S_{j}^{\prime}|}>\frac{m_{k}^{\prime}}{|S_{k}|}$ then
16:              $S_{j}^{\prime}=S_{j}^{\prime}\cup(S_{j}\cap S_{k})$
17:              $S_{k}^{\prime}=S_{k}^{\prime}-S_{j}^{\prime}$
18:           else
19:              break
20:           end if
21:        end for
22:     end if
23:  end for
24:  Return $\{G_{j}[0],S_{j}^{\prime}\},\forall j\in[1,n]$
@@ 4.2 Intersection Resource Scheduling (IRS)
We now tackle the first question, minimizing the scheduling delay. Directly matching jobs to devices can be computationally expensive, especially when dealing with the immense scale of devices and jobs. The challenges posed by this problem are not solely due to its scale; they are also compounded by the diverse resource requirements of CL jobs. Their various requirements introduce intricate resource contention patterns, further complicating the scheduling.
To this end, Venn introduces the Intersection Resource Scheduling (IRS) problem to account for this resource contention. Basically, each CL job $J_{i}\in\mathbb{J}$ may compete for a subset of devices $S_{k}\in\mathbb{S}$, denoted as $f(J_{i})=S_{k}$, where these resource subsets can exhibit relationships that are inclusive, overlapping, or nested. We created an integer linear programming (ILP) formulation (Appendix B) to optimally allocate resources to minimize scheduling delay and propose a heuristic to tackle the problem.
@@
With the objective of determining a job scheduling order, Venn first groups jobs $\mathbb{J}$ into Resource-Homogeneous Job Groups $\mathbb{G}=\{G_{1},G_{2},...,G_{n}\}$ by their resource requirements, where each job group $G_{j}=\{J_{i}|f(J_{i})=S_{j},\forall J_{i}\in\mathbb{J}\}_{i=1}^{m_{j}}$ contains all jobs with the same resource requirement. Venn addresses the problem using a two-step approach, with each step occurring at a different level of scheduling granularity, as outlined in Algorithm 1: (i) Ordering within a job group to optimize local resource scheduling (§4.2.1). (ii) Merging orders across job groups to ensure global scheduling efficiency (§4.2.2). Venn invokes Algorithm 1 on job’s request arrival and completion. By breaking down the overall problem into two steps, we further reduce the problem’s complexity without affecting the scheduling efficiency. We provide theoretical insights to show the effectiveness of this problem decomposition in Appendix C.
4.2.1 Intra-Job Group Scheduling
@@
4.2.2 Inter-Job Group Scheduling
@@
An effective scheduler should recognize and adapt to the resource contention patterns among jobs. Venn prioritizes groups with scarce resources (fewer eligible devices) to prevent delays from resource-rich groups. Additionally, when a particular resource type is in high demand, the scheduler should judiciously allocate intersected resources to the job group with a longer queue, as this group contributes more significantly to the average scheduling delay.
@@
    Amount of eligible resources allocated: the job group with smaller amount of eligible resources may have a longer scheduling delay under the same condition.
@@
Algorithm 1 outlines the steps Venn takes to allocate the current resource across job groups. First, Venn initializes resource allocation among job groups by starting to allocate resources to the job group with the most scarce resources (line 6). This results in an initial allocation plan with no resource sharing across job groups (lines 59), setting the stage for subsequent cross-group allocations.
To determine how to allocate intersected resources across job groups, Venn greedily evaluates whether a job group with more abundant resources should acquire intersected resources from groups demanding more scarce resources with the objective of minimizing the average scheduling delay. This evaluation starts with the job group possessing the most abundant resources (line 10). If the resources allocated to this group remain unclaimed by other groups (line 12), Venn will decide how much resources from subsequent job groups should be allocated to it. Specifically, Venn prioritizes job groups with longer queue lengths and fewer allocated resources, guided by a ratio that balances the number of affected jobs against the amount of allocated resources (line 15). If this ratio $\frac{m_{j}^{\prime}}{|S_{j}^{\prime}|}$ is larger than the one for the target resource group $\frac{m_{k}^{\prime}}{|S_{k}^{\prime}|}$, Venn reallocates resources accordingly (lines 1617). Otherwise, the algorithm ceases to allocate additional resources to the current job group from remaining groups (line 19). The reason is that if this job group needs more resources, it should first take the resources from job groups with relatively abundant resources.
Function get-queue-len() (line 14) would return the number of jobs $m^{\prime}$ whose JCT would be delayed by potential group prioritization. For example, the affected queue length $m^{\prime}$ may contain jobs from other job groups that have been deprioritized previously. An easier way to approximate $m^{\prime}$ is to use the group queue length. If intersected resources have been decided to be allocated to job group $G_{j}$ over $G_{k}$, Venn accumulates and updates their current resource allocation $S_{j}^{\prime},S_{k}^{\prime}$ and queuing length $m_{j}^{\prime}$ (line 1617).
The time complexity of Algorithm 1 is $\max(O(m\log m)$, $O(n^{2}))$, where $m$ is the number of ongoing jobs and $n$ is the number of job groups. We illustrate the theoretical insight behind the scheduling algorithm in Appendix D to illustrate the effectiveness of Venn’s approach.
@@ 4.3 Device Matching
Response collection time is usually determined by the last successfully responding devices. Hence, it can be reduced by allocating devices with similar higher capacity to the CL job. Meanwhile, these high-end devices have lower probability to fail due to their quick task execution.
However, as mentioned in Section 4.1, there is a trade-off between scheduling delay and response collection time. Intuitively, with limited device influx, priority should be given to minimizing the scheduling delay, which dominates the average JCT. On the other hand, with sufficient device influx to fulfill a job request within a short period, we should consider minimizing the response collection time while obeying the scheduling order given by Section 4.2.
@@
1:  Input: Jobs $J_{i}$, Devices $S_{j}^{\prime}\in{Venn}-SCHED(\mathbb{G},\mathbb{S})$
2:  $S=\{S^{1},S^{2},...,S^{V}\}$triangleright Evenly partition devices
3:  $g_{v}=\frac{t^{v}}{t^{0}},\forall v\in[1,V]$triangleright Response time speed-up
4:  Function: Venn-Match(Job $J_{i}$, Resource $S_{j}^{\prime}$)
5:  $c_{i}=\frac{t_{response}}{t_{schedule}}$triangleright Assign tiers in a rotating manner $\triangleright$ Decide whether to trigger tier-based matching
6:  $u=\texttt{randint}(0,V)$triangleright
7:  if $V+g_{u}\times c_{i}<c_{i}+1$ then
8:     Update $S_{j}^{\prime}=S_{j}^{\prime}\cap S^{u}$triangleright Assign tier $u$ devices to $J_{i}$
9:  end if
10:  Return: $\{J_{i},S_{j}^{\prime}\}$
@@
For each served job request, Venn randomly selects a device tier, denoted as $S^{u}$, to the job (Algorithm 2 line 5). This randomized tier assignment aims to expose each CL job to a diverse set of devices, rather than confining them to high-end devices. Given that the response collection time is determined by the slowest responding participant, tier-based assignment does not adversely affect this metric.
@@
To determine the response time speed-up factor $g$ for tier-based matching, we note that the device response time distribution adheres to a log-normal distribution (Wang et al., 2023). We use the 95th percentile as the statistical tail latency to account for the overall response collection time, thereby excluding failures and stragglers. Venn profiles and estimates the response collection time for each device tier $v\in[1,..,V]$ and subsequently computes the speed-up factor $g_{v}=\frac{t^{v}}{t^{0}}$ relative to a non-tiered scenario (line 3).
Figure 7: Visualized tier-based device-to-job matching condition.
@@ 4.4 Enhancements
Dynamic resource supply.
As shown in Figure 2(a), the total available CL resources change significantly over time. To address this, Venn continuously records each device eligibility through a time-series database. This database is then queried for resource eligibility distribution from the past time window. However, relying solely on momentary eligible resource rates for input into the scheduling algorithm is inaccurate. This is primarily due to the varying resource arrival patterns, as demonstrated in Figure 2(a). Since CL jobs span multiple days with diurnal resource patterns, Venn averages eligibility over 24 hours for robust scheduling. As a result, the scheduler can become both farsighted and robust, effectively accommodating the dynamic nature of resource availability.
Starvation prevention.
Our heuristic can lead to larger CL jobs being starved due to the preference given to smaller jobs. This is not acceptable especially when the jobs are initiated by different CL developers who require performance guarantees. Venn grants fairness to jobs to avoid such starvation. Specifically, our goal is to guarantee that the scheduling latency of a job $J_{i}$ is no worse than fair sharing, which is defined as $T_{i}=M*sd_{i}$, where $M$ is the number of simultaneous CL jobs and $sd_{i}$ represents the JCT without contention. Then, we adjust each job demand to be $d_{i}^{\prime}=d_{i}\times(\frac{t_{i}}{T_{i}})^{\epsilon}$ to ensure fairness within a job group, and adjust each group queue length $q_{j}^{\prime}=q_{j}\times(\frac{\sum_{J_{i}\in G_{j}}T_{i}}{\sum_{J_{i}\in G_{j}}t_{i}})^{\epsilon}$ to ensure fairness across job groups. $t_{i}$ is the time usage of job $J_{i}$ at the moment and $\epsilon\in[0,\infty)$ is a fairness control knob. When $\epsilon=0$, the algorithm is identical to the one in Section 4.2. As $\epsilon\rightarrow\infty$, the fairness multiplier dominates the scheduling, resulting in maximum fairness. We show that Venn improves JCT over its counterparts with our starvation design (§5.5).